# Unifews: Unified Entry-Wise Sparsification for Efficient Graph Neural Network

### v1

@@ 6.3. Efficiency on Decoupled GNN
(c) SGC on cora
(d) APPNP on cora
Figure 9. Sensitivity of joint sparsification thresholds on four models over cora. We respectively set the weight ratio $\delta_{w}$ to 00 and $0.5$, then evaluate the accuracy, edge sparsity, and weight sparsity of the model. Note that the x-axis representing diffusion threshold is on a logarithmic scale.
6.4. Ablation Study on Joint Simplification
[…]
Accuracy. We then showcase the ability of Unifews in conducting unified pruning on graph edges and network weights concurrently. Accuracy comparison is provided by Figure 6 with GCN backbone over cora as the example. Similar to Figure 4, the figures display the performance across varying edge and weight sparsity range, but with the other sparsity fixed at $\eta=30\%$. It is noticeable that Unifews retains comparable or better accuracy than the backbone GCN with up to $3\%$ improvement. Most baseline methods only obtain suboptimal accuracy especially in Figure 5(b), which is affected by their comparatively poor graph sparsification. The evaluation further highlights the advantage of Unifews in considering the two GNN operation stages in a unified manner.
[…]
FLOPs. The efficiency enhancement is explained by Figure 7, where we separately assess the FLOPs related to graph propagation and network transformation for different backbone models and datasets under Unifews sparsification. For iterative models, Figures 6(a) and 6(b) imply that the majority of computational overhead is network transformation, even on graphs as large as arxiv. Consequently, weight compression is essential for remarkably reducing GNN operations. On the other hand, graph propagation becomes the bottleneck in decoupled designs, and is increasingly significant on graphs with greater scales. This is because of the larger number of propagation hops $L=20$ for these structures. In this case, Unifews is effective in saving operational cost by simplifying the propagation and bypassing unnecessary computation, with over 20×20$\times$20 × and 5×5$\times$5 × joint reduction on arxiv and products, respectively. By combining these two sparsifications, we summarize that the unified scheme of Unifews is capable of mitigating the computational overhead for both propagation- and transformation-heavy tasks.
[…]
Sparsity. Additionally, we investigate the entry-wise sparsity of specific intermediate results in the process of GNN computation. Figure 8 displays the relative density of the edge message and node representation matrices, respectively. It can be clearly observed that the entry sparsity is enhanced with the increase of both two pruning ratios, signifying our theoretical analysis that Unifews not only directly shrinks edges and weights, but promotes the sparsity of the product matrix as well. It also supports our claim that Unifews is superior in employing dual sparsification where the two stages benefit each other alternatively and enjoy a win-win situation brought by the increased sparsity.
6.5. Effect of Hyperparameters
[…]
Sparsification Thresholds $\delta_{a}$ and $\delta_{w}$. We here perform a more in-depth analysis on the selection of the thresholds which are pivotal parameters in Unifews compression. Figure 9 denotes the changes of inference accuracy and dual sparsity of GNN models under graph and joint sparsification. Accuracy in the plot follows the conclusion in previous evaluation, that it only degrades above extreme sparsity $\eta_{a}>95\%$. The experiment reveals that, edge sparsity is positively related to the adjacency threshold $\delta_{a}$, and decoupled models require a larger range of threshold to traverse the sparsity range, which is because of the wider distribution of their entry values throughout the deeper propagation. Interestingly, the weight sparsity when $\delta_{w}=0.5$ also increases under high edge pruning ratios. The pattern is also observed in the reverse case. We deduce that the reciprocal enhancement in graph and model sparsity is brought by the unified sparsification of Unifews, which conforms to our analysis as well as previous studies (Zhang et al., 2023b).
[…]
Network and Propagation Layer $L$. Since Unifews adopts layer-dependent graphs with incremental sparsity for GNN propagation, we additionally evaluate its performance on models with varying layers. Figure 9(a) is the result of iteratively applying Unifews edge sparsification to SGC with different propagation depths. The backbone model without pruning $\delta_{a}=0$ typically suffers from the over-smoothing issue under large hop numbers. On the contrary, as elaborated in Section 5.2, Unifews is powerful for identifying and eliminating unimportant propagations, especially for larger hops, and thereby prevents information loss. With respect to architectural compression Figure 9(b), it is noticeable that Unifews promotes model performance and average weight sparsity at the same time for deeper layers, effectively reducing network redundancy.
@@ 6.5. Effect of Hyperparameters
Figure 10. Sensitivity of propagation and network transformation layers evaluated on SGC over cora. (a) Impact on prediction accuracy and average edge sparsity of varying numbers of propagation hops. (b) Impact on prediction accuracy and average weight sparsity of varying numbers of network layers. Note that the x-axis representing layers is on a logarithmic scale.
7. Conclusion
In this work, we present Unifews, an entry-wise GNN sparsification with a unified framework for graph edges and model weights. Unifews adaptively simplifies both graph propagation and feature transformation throughout the GNN pipeline and remarkably improves efficiency with little expense. In theory, we showcase that the layer-dependent Unifews provides an effective estimation on the graph learning process with a close optimization objective, which is favorable for multi-layer GNN updates. Comprehensive experiments underscore the superiority of Unifews in terms of efficacy and efficiency, including comparable or improved accuracy, $90-95\%$ operational reduction, and up to 100×100$\times$100 × faster computation.

### v2

@@ C.3 Effect of Hyperparameters
Figure 14: Accuracy of varying joint sparsification thresholds on GCN and GAT over cora.
@@
Figure 15: Sensitivity of propagation and network transformation layers evaluated on SGC over cora. (a) Impact on accuracy and average edge sparsity of varying numbers of propagation hops. (b) Impact on accuracy and average weight sparsity of varying numbers of network layers. Note that the x-axis representing layers is on a logarithmic scale.
Sparsification on Synthetic Graphs.
We further utilize GenCAT (Maekawa et al., 2022, 2023) to generate synthetic graphs with randomized connections and features by varying its parameters. GenCAT$(\alpha,\sigma)$ is the state-of-the-art graph generation approach that allows for configuring the edge connections and node attributes synthesis by specifying their distributions, where $\alpha$ is the coefficient of edge power law distribution $P(d)=d^{-\alpha}$, and $\sigma$ is the deviation of attribute Gaussian distribution $p[v]\sim N(0,\sigma^{2})$, both sharing the definitions in our paper. The remaining inputs of GenCAT, including label distribution and attribute-label correlation, are set by mimicking the statistics from cora. We utilize two groups of $\alpha$ and $\sigma$ values to generate 4 graphs with different edge and attribute distributions and evaluate Unifews.
Then, we evaluate the relationship between the edge threshold $\delta_{a}$ and the edge sparsity $\eta_{a}$ following the settings of Figure 8. The results are available in Figure 16. As an overview, the pattern is similar to the one in Figure 8, while the constants are effectively affected by the changes of edge and feature distribution in the GenCAT graph. For $\delta_{a}$ not too close to 00, the relation with $\eta_{a}$ largely follows Theorem 4.2.
Propagation and Network Layer $L$.
Since Unifews adopts layer-dependent graphs, we additionally evaluate its performance on models with varying layers. Figure 15(a) is the result of iteratively applying Unifews edge sparsification to SGC with different propagation depths. The backbone model without pruning $\delta_{a}=0$ typically suffers from the over-smoothing issue under large hop numbers. On the contrary, as elaborated in Section 4.2, Unifews is powerful for identifying and eliminating unimportant propagations, especially for larger hops, and thereby prevents information loss. With respect to architectural compression Figure 15(b), it is noticeable that Unifews promotes model performance and average weight sparsity at the same time for deeper layers, effectively reducing network redundancy.
C.4 Specific Experiments
Entry Distribution.
An example on the real-word graph cora is shown in Figure 9(a), which demonstrates that the distribution of node degree follows the power law, and the distribution of message values is correlated with the diffusion entries. Hence, message magnitude is effective in representing edge importance and determining entry removal. Setting entries in the diffusion matrix $\bm{T}$ to zero according to Equation 5 implies removing the corresponding edges from the graph diffusion process. Consequently, messages with small magnitudes are prevented from propagating to neighboring nodes and composing the output embedding.
Approximation Error.
An empirical evaluation of multi-hop Unifews on cora is shown in Figure 9(b), which validates that the approximation error is affected by the sparsification threshold $\delta_{a}$ as well as propagation hops $L$. Meanwhile, even under aggressive sparsification, the error is still adequately bounded to yield meaningful learning outcomes.
Over-smoothing.
Additionally, we note that the “error” induced in Algorithm 1 is not necessarily harmful. Examining Equation 1, excessive propagation of common GNNs can lead to a decline in performance known as the over-smoothing issue, where the graph regularization is dominant and meaningful node identity is lost (Li et al., 2018). By eliminating a portion of graph diffusion, Unifews effectively alleviates the over-smoothing issue.
We showcase the effect of alleviating over-smoothing in Unifews by depicting the embedding difference to optimization convergence in Figure 9(c). With an increased number of hops, the output embedding tends towards an over-smoothed state. However, a stronger sparsification prevents the rapid smoothing and hereby contributes to better graph learning performance.
We further study the particular effect with the iterative GCNII of 32 layers in Figure 12, and with the decoupled SGC of 5-80 layers in Figure 15. Both results demonstrate that, by increasing model layers, the accuracy of the pruned model is largely preserved thanks to the residual connections. Hence, we conclude that Unifews pruning also benefits accuracy by alleviating over-smoothing.
(a) GenCAT$(2.1,0.1)$
(b) GenCAT$(2.1,1.0)$
(c) $\alpha=2.1$ Degree Distribution
(d) GenCAT$(2.9,0.1)$
(e) GenCAT$(2.9,1.0)$
(f) $\alpha=2.9$ Degree Distribution
Figure 16: Relationship between edge threshold $\delta_{a}$ and sparsity $\eta_{a}$ on synthetic graphs generated by GenCAT. We also present the degree distribution of nodes in different GenCAT graphs controlled by $\alpha$ to illustrate the power law.